### 12 Dec 2023 v1

Abstract
[…]
keywords:
isospin symmetry breaking, QCD, kaon production
In 1932, Heisenberg suggested that the neutron and the proton could be seen as different manifestations of the same particle, the nucleon [1], explaining their similar masses. Wigner referred to this feature as ‘isotopic spin’, or ‘isospin’, due to the formal similarity with the more familiar spin [2]. Today, isospin is an important element of nuclear physics and nuclei classification [3]. The concept of isospin was promptly extended to the other strongly interacting particles, the hadrons. In particular, Kemmer [4] applied it to the pion triplet postulated by Yukawa in 1935 [5]. For a historical outline, see Refs. [6, 7].
Nowadays, hadrons are viewed as emergent within Quantum Chromodynamics (QCD) [8], the theory of strong interactions between the elementary carriers of colour charge, the quarks and the gluons. Because of confinement, QCD is based on a local colour gauge symmetry that is hidden in colourless hadrons. Hadrons are classified into mesons with integer spin, typically quark-anti-quark states such as pions and kaons, and baryons with semi-integer spin, typically three-quark states such as protons and neutrons.
Within QCD, the isospin symmetry could be understood as a part of a more general symmetry, denoted as flavour symmetry, which arises because the gluons interact ‘democratically’ with the light quark flavours $u$, $d$, and $s$. Moreover, in the limit where the quarks can be considered massless (the so-called chiral limit), these interactions do not depend on their chirality, thus realizing a larger symmetry called chiral symmetry. The latter is spontaneously broken in the QCD vacuum generating (quasi-) Goldstone bosons, pions, kaons, and the $\eta$-meson.
In more specific terms, the isospin operator triplet $\hat{\mathbb{I}}=(\hat{I}_{x},\hat{I}_{y},\hat{I}_{z})$ refers to an $SU(2)$ rotation of the ‘isodoublet’ $(u,d)$, where $u$ has eigenvalue $I_{z}=1/2$ and $d$ $I_{z}=-1/2$, thus an isospin transformation changes e.g. a proton into a neutron or generates an $O(3)$ rotation of the pion triplet. Kaons, which are grouped into the isodoublet $(I=1/2)$ states $(K^{+}\equiv u\bar{s},K^{0}\equiv d\bar{s})$ and $(\bar{K^{0}}\equiv s\bar{d},K^{-}\equiv s\bar{u})$ correspond naturally to the underlying isospin doublets $(u,d)$ and $(-\bar{d},\bar{u})$ and shall play a crucial role in the following. Isospin symmetry implies that only transitions preserving total isospin $I$ and its $z$-component $I_{z}$ are allowed. This property is referred to as the isospin conservation.
QCD would strictly obey the isospin symmetry, having equal light quark masses, $m_{u}=m_{d}$. Since the current quark masses of $u$ and $d$ quarks differ only by a few MeV [9], only small deviations from this symmetry are expected. The relative mass differences of charge states of pions and kaons are small, $\left(m_{\pi^{+}}-m_{\pi^{0}}\right)/\left(m_{\pi^{+}}+m_{\pi^{0}}\right)\simeq 0.017$ and $\left(m_{K^{+}}-m_{K^{0}}\right)/\left(m_{K^{+}}+m_{K^{0}}\right)\simeq-0.004$. The elastic cross sections for pion-pion, pion-nucleon, and nucleon-nucleon scattering closely follow the predictions of isospin symmetry [10, 11].
Thus, one also expects that isospin symmetry is obeyed in good approximation in hadron production processes at high energies. However, recently reported NA61/SHINE experimental results on producing charged and neutral kaons, supported by the compilation of the world data reported in Ref.  [12], seem to contradict those expectations [13]. Here, we address this question in detail.
Isospin symmetry and kaon production in nucleus-nucleus collisions.
Let us consider, for definiteness, an initial state with two identical nuclei having equal proton and neutron numbers ($Z=N=A/2$) in their ground states [^1] Considering two distinct nuclei, each of them having an equal number of protons and neutrons, would lead to the same results.. This initial state is denoted as $\left|A+A\right\rangle$. In conserved quantities, the condition $Z=N$ is given by $Q/B=1/2,$ where $Q=Z$ and $B=A$ are the nuclei’s electric charge and baryon number, respectively. Then, each nucleus’ third component of isospin is zero [^2] For almost all nuclei of this type, the total isospin is also zero [14], hence the total isospin of the $\left|A+A\right\rangle$ state vanishes..
The isospin transformation is parameterized by the unitary operator $\hat{O}=e^{i\theta_{j}\hat{I}_{j}}$ and, in general, (approximate) isospin symmetry of QCD means that $\hat{O}^{\dagger}\hat{H}_{QCD}\hat{O}\simeq\hat{H}_{QCD}$, where $\hat{H}_{QCD}$ refers to the Hamiltonian of QCD. A particularly useful unitary operator is the special isospin transformation, called charge symmetry transformation, $\hat{C}=e^{i\pi\hat{I}_{y}}$. It corresponds to an inversion of the third component of isospin: $\hat{C}\left|u\right\rangle=-\left|d\right\rangle$ and $\hat{C}\left|d\right\rangle=\left|u\right\rangle$, schematically $u\leftrightarrow d$. At the hadronic level, the $\hat{C}$ transformation implies $p\leftrightarrow n$, $\pi^{+}\leftrightarrow\pi^{-}$, $K^{+}\leftrightarrow K^{0}$, $\bar{K}^{0}\leftrightarrow K^{-}$, etc.
Without referring to a detailed mathematical formalism, charge symmetry means that strong interactions are invariant under the inversion of the third component of the isospin of every nucleus and hadron of the initial and final states. Let us consider an ensemble of initial states being invariant under the charge transformation $\hat{C}$ probabilities of having initial states related by this transformation are equal. This is indeed the case of nucleus-nucleus collisions where each nucleus has an equal number of protons and neutrons (thus, $I_{z}=0$). Then, the invariance under $\hat{C}$-transformation holds also for the final state ensemble, implying that the mean multiplicities of $\hat{C}$-related quantities, such $K^{+}$ and $K^{0}$ and $\bar{K}^{0}$ and $K^{-}$, coincide:
$\langle K^{+}\rangle=\langle K^{0}\rangle\text{ , }$ (1)
and
$\langle K^{-}\rangle=\langle\bar{K}^{0}\rangle\text{ . }$ (2)
Note, since models predict only properties of ensembles of events, we need to consider mean multiplicities for testing charge invariance.
The relations (1) and (2) are in agreement with the result first formulated in 1955 by Shmushkevich [15, 16], see also Refs. [17, 18, 19]. This result states that an initial ‘uniform’ ensemble of hadronic state (that is, one with an equal mean number of each member of any isospin multiplet) evolves into a uniform final-state ensemble.
The neutral kaons are eigenstates of strong interaction, oscillating ($K^{0}\leftrightarrow\bar{K}^{0}$) due to weak interaction. The weak eigenstates are the long-lived ‘K-long’ $K_{L}^{0}$ and short-lived ‘K-short’ $K_{S}^{0}$, which decay into three and two pions with a mean lifetimes of 15 m/$c$ and2.8 cm/$c$, respectively. By neglecting the small CP violation, the weak and strong kaonic eigenstates are related by:
$\left(\begin{array}[c]{c}\left|K_{S}^{0}\right\rangle\\ \left|K_{L}^{0}\right\rangle\end{array}\right)=\frac{1}{\sqrt{2}}\left(\begin{array}[c]{cc}1&-1\\ 1&1\end{array}\right)\left(\begin{array}[c]{c}\left|K^{0}\right\rangle\\ \left|\bar{K}^{0}\right\rangle\end{array}\right)\text{ ,}$ (3)
hence, for the mean number of $K_{S}^{0}$ and $K_{L}^{0}$ one gets:
$\langle K_{S}^{0}\rangle=\frac{1}{2}\langle K^{0}\rangle+\frac{1}{2}\langle\bar{K}^{0}\rangle=\langle K_{L}^{0}\rangle\text{ .}$ (4)
Thus, we have an equal number of long-lived and short-lived kaons, even if $\langle K^{0}\rangle$ and $\langle\bar{K}^{0}\rangle$ are different due to the non-zero baryon number of the initial state.
By combining Eqs. (1) and (2) and using Eq. (4) one finds:
$\langle K^{+}\rangle+\langle K^{-}\rangle=2\langle K^{0}_{S}\rangle~{},$ (5)
where also the r.h.s. in accessible to experiments. Consequently, the exact isospin symmetry prediction for the charged-to-neutral kaon ratio in nucleus-nucleus collisions with $Q/B=1/2$ reads:
$R_{K}\equiv\frac{\langle K^{+}\rangle+\langle K^{-}\rangle}{\langle K^{0}\rangle+\langle\bar{K}^{0}\rangle}=\frac{\langle K^{+}\rangle+\langle K^{-}\rangle}{2\langle K_{S}^{0}\rangle}=1~{}.$ (6)
This prediction serves as the initial reference for testing the isospin symmetry with the results provided by the NA61/SHINE experiment [12] that measured $R_{K}=1.262\pm 0.017(stat)\pm 0.062(sys)$ in collisions of argon and scandium nuclei at the energy of 11.9 GeV per nucleon pair. This result is consistent with the compilation of the world data, even if each previous measurement, taken individually, was also consistent with $R_{K}=1$ within its large uncertainty. Yet, as discussed below, all data taken together clearly show a significant deviation from unity.
Neither is the isospin symmetry exact nor does experimental data refer to $A+A$ collisions with $Q/B=1/2$. In the following, we estimate possible deviations of $R_{K}$ from unity due to the known effects: isospin symmetry-breaking processes and $Q/B\neq 1/2$, using popular and well-established statistical and dynamical models. Then, we quantify the difference between the experimental results and the model predictions and discuss possible implications.
Known isospin symmetry-breaking effects. The isospin symmetry is explicitly broken in weak and electromagnetic processes and is an approximate symmetry of strong interactions. Here, we consider the known isospin symmetry-breaking effects one by one.
  1. (A)
    Weak processes. The weak interaction does not obey the isospin symmetry. The mean lifetimes [9] of charged and neutral kaons are $\tau(K^{+})=\tau(K^{-})=(1.238\pm 0.001)\cdot 10^{-8}$ s and $\tau(K^{0}_{S})=(8.954\pm 0.004)\cdot 10^{-11}$ s, $\tau(K^{0}_{L})=(5.116\pm 0.021)\cdot 10^{-8}$ s. The charged kaons are typically measured by reconstructing their trajectories in a detector. Due to the large mean lifetime, the corrections for their losses caused by weak decays are small. In contrast, the neutral $K^{0}_{S}$ kaons are measured by reconstructing their decays into two charged pions. Typically, the corrections for the losses caused by weak decays are large. This is because the decay should be far enough from the interaction point to separate the decay point from the background in high multiplicity $A+A$ collisions. Assuming that the $K^{0}_{S}$ mesons are measurable when their proper lifetime is larger than the mean lifetime (typical for NA61/SHINE [12]), one estimates the maximum relative bias of the mean multiplicity as
    $\frac{\Delta(\langle K_{S}^{0}\rangle)}{\langle K_{S}^{0}\rangle}=\frac{3\cdot\sigma(\tau(K_{S}^{0})}{\tau(K_{S}^{0})}\approx 0.0013.$ (7)
    Thus, the maximum deviation of $R_{K}$ from unity due to the uncertainty of the mean $K^{0}_{S}$ lifetime is 0.13 %.
  2. (B)
    Electromagnetic process. Electromagnetic interactions do not conserve isospin because the electromagnetic coupling constant differs for the quark flavours $u$ and $d$. Yet, electromagnetic decays of hadrons are typically suppressed by a factor $\alpha\simeq 1/137$ compared to strong ones. Consequently, processes that involve the production of virtual photons and their subsequent decay into charged kaons are suppressed by a factor $\alpha^{2}$ and thus negligible.
  3. (C)
    Strong processes within QCD. The isospin symmetry is not exact because $u$ and $d$ quark masses are different, $\approx 2.2$ and $\approx 4.7$ MeV, respectively. The different quark masses lead to different masses of hadrons within the isospin multiplets, particularly different masses of charged and neutral kaons. The ratio $R_{K}$ can be expanded as a function of the mass difference of the $u$ and $d$ quarks as:
    $R_{K}=R_{K}^{(0)}+\frac{m_{d}-m_{u}}{\Lambda_{QCD}}R_{K}^{(1)}+...\text{ ,}$ (8)
    where $R_{K}^{(0)}$ is the value of the ratio for the isospin symmetry ($m_{d}=m_{u}$) being equal to one for $Q/B=1/2$, and $\Lambda_{QCD}\simeq 250$ MeV is the QCD energy scale [20, 21]. The question is whether $R_{K}^{(1)}+...$ terms can reproduce the experimental data. We list below the considered effects and quantify their impact on $R_{K}$ using the statistical hadron resonance gas model (HRG) [22]. The results are cross-checked with the microscopic, UrQMD transport model [23, 24, 25]. For details, see Methods.
    1. (i)
      Smaller masses of charged kaons than neutral ones, $m_{K^{+}}=m_{K^{-}}=493.7$ MeV and $m_{K^{0}}=m_{\bar{K}^{0}}=497.6$ MeV, lead to an increase of $R_{K}$ resulting from direct kaon production by about 0.02. This was estimated by removing resonances from the particle list of HRG.
    2. (ii)
      Different kaon masses impact the branching ratios of hadronic resonances decaying into kaons. The most striking example is $\phi(1024)$ meson, which decays twice as frequently to charged kaons as into neutral ones. This large difference is because the $\phi(1024)$ mass is just above the kaon-kaon thresholds. Including the kaon production from resonance decays increases $R_{K}$ by about 0.01.
    3. (iii)
      Mass differences of hadrons belonging to other isospin multiplets also break the isospin symmetry and affect $R_{K}$. The largest effect comes from the mass difference between proton and neutron, which reduces $R_{K}$ at the lowest collision energies. At energies larger than 10 GeV, this effect is negligible.
    Interestingly, HRG for $Q/B=1/2$ and in the exact isospin-symmetry limit reproduces Eq. (1) exactly.
Figure 1: The charged-to-neutral kaon ratio $R_{K}$ as a function of collision energy. The symbols show the experimental data. The black line shows the HRG predictions for $Q/B=0.4$. The blue dots indicate the HRG predictions for $Q/B$ values corresponding to the ones in experiments. For different nuclei, $Q/B$ corresponds to the electric charge over baryon number of the whole system. The grey rectangles show the UrQMD model results.
Collisions with $Q/B<1/2$. For nuclei with $Q/B<1/2$ the third component of isospin equals $|I_{z}|=|B/2-Q|$ and therefore the total isospin is limited as $|I_{z}|\leq I\leq B/2$.
The compiled experimental results in nucleus-nucleus collisions correspond to the $Q/B$ ranging from about 0.4 (Pb+Pb and Au+Au collisions) to 0.5 (Si+Si and S+S collisions); see Fig. 1. The limits correspond to the normalized per baryon third component and total isospin $|I_{z}|/B=0.1$, $20.1<I/B<1/$ 2 and $|I_{z}|/B=I/B=0$, respectively. The non-zero $I_{z}$ and $I$ for heavier nuclei can impact the charged-to-neutral kaon ratio in two ways:
  1. (a)
    The larger fraction of neutrons than protons for heavy nuclei having $|I_{z}|/B\approx 0.1$ enhances neutral kaon production compared to charged ones and thus reduces $R_{K}$. This increases the disagreement between the experimental results and the model predictions calculated for $|I_{z}|=I=0$. The reduction is significant at low collision energies and is small compared to other effects at energies larger than 10 GeV; see Fig. 1 and Methods.
  2. (b)
    For $|I_{z}|/B\approx 0.1$, the total isospin is limited by $20.1\leq I/B\leq 1/$ 2. Generally, nuclei in the ground state have the lowest possible value of the total isospin [14]. This rule extends to a state of two identical nuclei in the ground state, which, for the considered case, implies $I/B\approx|I_{z}|/B\approx 0.1$. Thus, in the considered cases, a possible dependence of $R_{K}$ on $I$ and $I_{z}$ reduces to the dependence on $I=|I_{z}|$. The latter is discussed in (a) above.
    The experimental data also allow a rough estimate of the influence of the normalized isospin small but non-zero value on the charged-to-neutral kaon ratio. Results for of heavy nuclei, ${}^{208}$b+${}^{207}$b and ${}^{197}$u+${}^{197}$u ($R_{K}\approx 1.15$ for $I/B=|I_{z}|/B\approx 0.1$), are similar to those for intermediate nuclei, ${}^{40}$r+${}^{45}$c ($I/B=|I_{z}|/B\approx 0.08$), and ${}^{28}$i+${}^{28}$i ($I/B=|I_{z}|/B=0$), see Fig. 1. This suggests that the sensitivity of the ratio to the $I/B=|I_{z}|/B$ close to zero ($I/B=|I_{z}|/B\leq 0.1$) is low. Note, however, that there are large uncertainties in the experimental results.
In the case of central collisions, the charge-to-baryon ratio of the interacting nucleon system (participant nucleons) can be higher than the total proton-to-nucleon ratio in colliding nuclei. This is because protons tend to be distributed closer to the centre of a nucleus. However, this effect is expected to be rather small and relevant only for the lowest energies ($\lesssim 4$ GeV), where $R_{K}$ is sensitive to the small changes of $Q/B$.
In summary, at energies larger than 10 GeV, the mass difference between charged and neutral kaons, implying isospin symmetry-breaking processes, increases the $R_{K}$ ratio by about 0.03. Other isospin-breaking effects can be neglected. The ratio $R_{K}$ is reduced for collisions with $Q/B<1/2$, but this is insignificant for energies larger than 10 GeV. At energies lower than 10 GeV, the $R_{K}$ ratio is significantly more sensitive to the known isospin-breaking effects and the $Q/B$ ratio; see Methods for more details.
The above findings are illustrated in Fig. 1 showing predictions of the HRG and UrQMD models as functions of collision energy. Importantly, the plot includes the experimental results concerning the $R_{K}$ ratio from NA61/SHINE [12] and from other experiments, namely CERES [26], STAR BES [27, 28], STAR [29, 30, 31, 32], ALICE [33, 34], NA35 [35, 36, 37], NA49 [38], HADES [39, 40, 41, 42] and FOPI [43, 44]. The compilation is taken from the NA61/SHINE publication [12]. The HRG calculations were performed with $Q/B=0.4$ (solid line) and $Q/B$ values corresponding to collisions studied in the experiments (black dots). The UrQMD results were obtained for central Au+Au collisions. See Methods for details.
The two main observations are that the model predictions agree well with each other but are systematically lower than the experimental data.
Figure 2: The experimental data for the charged to neutral kaon ratio divided by the HRG baseline $R_{K}/R_{K}^{\rm HRG}$ as a function of collision energy. The solid black line shows the weighted average of the experimental data, and the shaded area shows the error of the weighted average.
Quantification of the isospin-symmetry breaking beyond known effects. To quantify the measured isospin-symmetry breaking (in short, isospin-breaking) beyond the known effects, the ratio of the measured $R_{K}$ to the corresponding HRG one is shown in Fig. 2. We do not consider the three lowest energy data points (HADES and FOPI) because the known isospin-breaking and $Q/B<1/2$ effects are significant at the low collision energies. Thus, the predictions may be model-dependent. We also do not consider two NA35 points because, unlike other measurements, charged kaons were identified by reconstructing their decays, leading to large statistical uncertainties and possible biases. Thus, the number of selected measurements at different collision energies (from AGS to LHC) is $15$. Out of them, only one is below unity. No significant dependence of the double ratio on collision energy and nuclear mass number of colliding nuclei is visible.
The weighted average of all results is $1.143\pm 0.026$, where the second number stands for the error calculated following the Particle Data Group procedure [9]. The weighted average deviates from the expectation for the isospin symmetry broken by the known effects by $5.55$ errors. The $\chi_{\rm min}^{2}/\text{dof}=0.47$ may indicate an overestimation of the uncertainties. The estimated scaling factor for the uncertainty is 2.15, and the corresponding scaled error is 0.018. The weighted average deviates from the expectation for the isospin symmetry, broken by the known effects by 8.14 scaled errors.
Closing remarks. Several comments are in order here.
  1. (I)
    The known effects of the isospin-symmetry breaking fail to explain the experimental results on the charged-to-neutral kaon ratio reported recently by NA61/SHINE [12]. A dedicated experimental and theoretical effort is needed to solve the puzzle of this unexpected large isospin symmetry breaking.
  2. (II)
    Concerning measurements, reviewing the validity of the past results and confirming them with new high-precision data is important. Systematic results on the collision energy and nuclear mass dependence of the isospin-breaking effect should help us understand its nature. Measurements of the charged-to-neutral kaon ratio in collisions of an equal number of protons and neutrons would eliminate uncertainty in its interpretation. In particular, measurements of the ratio in deuterium-deuterium interactions, the lightest nuclei with $I=0$, are important.
  3. (III)
    An interesting experimental test is also possible by considering $\pi^{-}+C$ and $\pi^{+}+C$ interactions. While the ensemble of only one of them is not invariant under charge transformation, the ensemble having an equal number of $\pi^{-}+C$ and $\pi^{+}+C$ interactions is invariant. Thus, for the joint ensemble, the exact charge symmtery predicts $R_{K}=1$.
  4. (IV)
    Kaons play a special role due to the simple isospin structure and easy-to-measure measure. This explains why the first results on a large isospin-symmetry breaking are reported for kaons. Yet, it is important to perform a similar study for other isospin multiplets in the future. For example, using the same methods, we have checked that the (anti-)proton to (anti)-neutron ratio is even less sensitive to the known isospin symmetry-breaking effects and, thus, is predicted to be almost exactly one in nucleus-nucleus collisions with $Q/B=0.5$ in a broad range of collision energies including the low energies. Considering collision-by-collision fluctuations of charged and neutral kaon numbers would also be interesting. A puzzling result was recently reported by ALICE at LHC [45] and discussed in Ref. [46].
  5. (V)
    The relationship between the production rates of charged and neutral kaons in hadronic collisions was used to estimate the flux of secondary neutral kaons produced in $p$+Be collisions at beam momenta of several 100 GeV/$c$ in neutral kaon and neutrino beams [47]. Recently, the charged-to-neutral kaon ratio was studied in p+p interactions at SPS energies, and no excess in comparison to a simple quark counting model emerged [48]. On the other hand, $\pi^{-}$+C interactions at SPS energies yield the ratio of charged-to-neutral kaons, which is significantly larger than unity and predictions of popular models [49].
  6. (VI)
    Concerning models, one may argue that there are known isospin-breaking effects which are improperly implemented in the popular models. For example, an isospin-breaking in production and decay of heavy resonances could sum up a significant result. While this seems hard to hold since we considered known effects of this type, careful and independent verification is required. Independently, one can extend the models by introducing new isospin-breaking processes and fitting their parameters to the data. This can be done either for the quark-gluon processes or the hadron-resonance processes. For example, within the statistical hadronization models, one can introduce the quark fugacity factors for $u$ and $d$ quarks separately [50, 51, 52]. This could allow us to make predictions for other hadron ratios.
  7. (VII)
    Concerning unknown effects violating isospin symmetry, the possibility of having a phase of strongly interacting matter with a significant isospin violation was suggested by Pisarski and Wilczek within a linear $\sigma$ model of QCD [53]. They expect masses of $\pi^{0}$ and $\eta$ mesons to decrease if the $U_{A}(1)$ symmetry is effectively restored at lower temperatures than the chiral phase transition. Can the chiral anomaly [54, 55] or other non-perturbative phenomena within QCD significantly affect the charge-to-neutral kaon ratio?
[…]
The evidence for unexpected large isospin symmetry breaking for kaons that cannot be explained by otherwise successful models seems well upheld. Assuming it passes future experimental tests, the quest is to understand its origin.
[…]
Acknowledgments
The authors thank Wojciech Broniowski, Katarzyna Grebieszkow, Stanislaw Mrowczynski, Owe Philippsen, Rob Pisarski, Krishna Rajagopal, Jan Steinheimer, Leonardo Tinti, Ludwik Turko and Volodymyr Vovchenko for fruitful discussions and comments. This work is partially supported by the Polish National Science Centre (NCN) grants 2018/30/A/ST2/00226 and 2019/33/B/ST2/00613, the German Research Foundation grants GA1480/8-1 and the Alexander von Humboldt Foundation.
Methods
Hadron Resonance Gas Model
We use the Hadron Resonance Gas model implementation from Ref. [22] to quantify the isospin-breaking effects and their interplay. HRG includes all hadrons and resonances with confirmed status in the Particle Data Tables [56]. The PDG-listed masses, charges, lifetimes, and decay modes are used. Thus, HRG includes the isospin-symmetry violation due to masses and branching ratios of hadrons and resonances.
In HRG calculations, the exact net strangeness conservation is enforced, i.e., the calculations are done within the strangeness canonical ensemble (SCE) [57, 58]. The model parameters are baryo-chemical potential, $\mu_{B}$, temperature, $T$, volume of the system $V$, and the strangeness under-saturation parameter $\gamma_{S}$ (see Ref. [59]). We adopt a simple parametrization of $\mu_{B}$ and $T$ as a function of collision energy introduced in Ref. [60]. We have checked that $R_{K}$ is weakly sensitive to the strangeness suppression effect introduced by having the parameter $\gamma_{S}<1$. Therefore, for simplicity, calculations are done for $\gamma_{S}=1$. The energy-dependent Breit-Wigner spectra [61] model the resonance widths.
Figure 1 (black line) shows the HRG predictions for $R_{K}$ as a function of collision energy $\sqrt{s_{NN}}$ for $Q/B=0.4$. At low collision energies, $R_{K}<1$ due to the enhancement of neutral kaon production caused by a larger number of neutrons than protons, $Q/B<1/2$. On the other hand, at high collision energies, HRG predicts $R_{K}\approx 1.018$ due to the mass difference between charged and neutral kaons produced directly. Finally, $R_{K}\approx 1.032$ when kaon production via resonance decays is included. This is mostly due to $\phi$ decays, which strongly prefer the decay into charged kaons over the one into neutral kaons.
We have checked that the system’s electric to baryon charge ratio $Q/B$ significantly affects $R_{K}$ up to $\sqrt{s_{NN}}\approx 10$ GeV. At higher energies, pions dominate, and total electric and baryon charges are significantly larger than the corresponding net charges. Thus, $R_{K}$ becomes increasingly less sensitive to $Q/B$ with an increase in $\sqrt{s_{NN}}$.
We have checked that the strangeness grand-canonical ensemble and other popular parametrizations [62, 63] of the model parameters as a function of collision energy lead to quantitatively similar results for $R_{K}$, for $\sqrt{s_{NN}}\gtrsim 4$ GeV. The uncertainties of $R_{K}$ estimated by changing the parametrization of model parameters [60] are less than $1$% for $\sqrt{s_{NN}}>3$ GeV. The effect of including light nuclei in the particle list is negligible.
UrQMD model
The UrQMD transport model [23, 24, 25] describes A+A collisions by explicitly propagating hadrons in phase space. During the propagation, rescattering among hadrons takes place. The particle production in this model happens either via resonance decay or via a string excitation and fragmentation following the LUND model [64]. The grey rectangles in Fig. 1 indicate the UrQMD model predictions. Here, we have considered central Au+Au collisions ($A=197,Z=79$, $Q/B\approx 0.4$). The predictions are shown within the $\sqrt{s_{NN}}$ range of 2.4 to 20 GeV. At each energy, $10^{4}$ events are used for the analysis.
One sees that for $\sqrt{s_{NN}}\lesssim 7$ GeV, the predictions of UrQMD and HRG are similar. At higher energies, the $R_{K}$ ratio in HRG is systematically higher than the one predicted by UrQMD. This is likely caused by UrQMD assuming $\phi$-meson decays to be exactly isospin symmetric instead of taking the branching ratios for PDG. This is the reason for showing the UrQMD predictions only up to 20 GeV.